A monopolist demand function is given by; P=Q2-10Q+28 and cost function; C=Q2. If the firm is causing a pollution to the society valued at MC=3, solve for the social optimal level of output and price. What should be the tax level?
1
Expert's answer
2020-05-21T11:17:41-0400

The social optimal quantity is estimated by equating the price to be equal to the marginal cost.

"\text{P=MC}"

Therefore:

"3=Q^2-10Q+28"

"Q^2-10Q+28-3"

"Q^2-10Q+25"


"\dfrac {-b \pm \sqrt{b^2-4ac}}{2a}"


"\dfrac {10 \pm \sqrt{ (-10)^2-4\times 1 \times 25}}{2\times 1} =5"


The social optimal level is therefore, 5 units.


The optimal price is:

"p=5^2-10\times 5+28"

"p=3"

Or

"\text{P=MC}"

"\text{P=3}"


At social optimum level, there will be no taxation instead the government should instead give subsidies to the monopolist.